wat is the value of 1/0?
 

Grade:12th pass

22 Answers

Mondo Camry
8 years ago
it is infinity
Prajwal Kavad
8 years ago
infinity meaaning
Adarsh
8 years ago
1/0 cannot be defined or its value can be taken as infinity
deapan
8 years ago
answer is infinite
noogler
8 years ago
the value of 1/0 is undefined bcuz
if we divide 1 with 0 
     0)1(123456789                              0)1(9999999
        0                                                 0
        0                                                  0
        .                                                   .
        .                                                   . 
        .                                                   .
 
in the same way we can use diff numbers.so its value cannot be defined
utpal raj
7 years ago
we take 0 as 0.0000000011 some number and divide ,
so 1/0.0000000011=909090909.1 
which is an infinite number  , so ans is infinity
 
Lakshmipathi
7 years ago
It is not infinity.It is undefined because the meaning of the both are different.
REDDY PAVAN
7 years ago
it is infinite value because any thing by zero is not defined and hence considered as infinte value......
jyoti bhatia
7 years ago
If we multiply 1/0, we could get 0 or 1.
We cannot define 1/0 to be a number. We cannot have both possibilities so we cannot define 1/0 to be a number. So, it is undefined.
Ritika Das
6 years ago
It is generally not at all possible to divide one by zero. Instead we call it as undefined. But the scientific explanation for it is that it is equal to infinity.
Agastya Varma
6 years ago
1/0 is actually not defined in fact any real number divided by zero is not defined. But the the value of 1/0 is very close to infinity but not infinity.
Agastya Varma
6 years ago
@prajwal infinity is defined as a very large number that has no known value. It is represented by the symbol \infty.
Nitesh
6 years ago
Answer is infinity. Because 0 cannot divide 1 at the infinite times.If we divide 1 by 0 it goes to infinite time multiple of 0 with 0 and after that also we cannot found result.So it is said that the answer is infinity.
Harshit Panwar
6 years ago
1/0 is not defined in maths point of view whereas it is infinity from physics point of view..........
Hope it helps you..............
Kochouseph
6 years ago
Really 1/0 is not defined. But we study in smaller classes that 1/0=0 because if the denominator is 0 then the value is 0
Priyansh
6 years ago
infinity as zero multiplied by anything gives zero thus division by zero is not defined...........................................................................
Om
5 years ago
Dear student
Any number divided by zero is not defined
i.e 1/0 is not defined 
I hope this answer helps you
With regards 
A fellow student
Soma Mesh
5 years ago
The answer is very simple it is infinity because any no divisible by zero is actually infinity.it can be explained simply that it can't be divided.
john cena
5 years ago
dear student
every rational number should be in the form of p/q where q is not equal to zero
but here
the value of denominator cannot be zero
so it determines that 1/0= infinty (\infty)
Bharat Luthra
5 years ago
The value of 1/0 is not defined or is infinity, because there is no number with which we multiply 0 to get1.
Hope it helps..
REENU SIVADARSHINI.M
5 years ago
1\0 = INFINITY OR UNDEFINED SO ANY NUMBER DIVIDED BY ZERO IS UNDEFINED.
**************************************************************************************************
Sayantan Garai
5 years ago
We take any number divided by 0 as undefined.
If the numerator is not 0 then division by 0 is taken as infinity.
Example
n/0 is undefined, if n=/= 0 ( n not equal to 0) then n/0 is infinity.